On July 1, 2019, Sandhill Co. purchased new equipment for $90,000. Its estimated useful life was 8 years with a $18,000 salvage value. On December 31, 2022, the company estimated that the equipment’s remaining useful life was 10 years, with a revised salvage value of $5,000.

Compute the revised annual depreciation on December 31, 2022.

Revised annual depreciation

$

Could you please show working?
Answer #1

Answer

  • Revised annual Depreciation = $ 6250 Answer
  • Workings

A

Original cost

$                     90,000

B

Original salvage value

$                     18,000

C = A - B

Original depreciable base

$                     72,000

D

Original expected life [years]

8

E = C/D

Annual 12 month depreciation

$                       9,000

F = E x 2.5 years

Total Accumulated Depreciation till estimates changed for 2.5 years [6 months 2019 + 12 months 2020 + 12 months of 2021]

$                     22,500

G = A - F

Book Value at the time of change of estimate

$                     67,500

H

New salvage value

$                       5,000

I = G - H

New depreciable base

$                     62,500

L

Remaining life

10

M = I/L

Revised depreciation expense from 2022

$                       6,250 ANSWER